OpenStudy (anonymous):

At what rate is the diagonal of a cube increasing if its edges are increasing at a rate of 2 cm/s?

OpenStudy (anonymous):

the diagonal of a cube = \[d = \sqrt{3x^2}=\sqrt{3}x\] so the length of the diagonal of a cube is proportional to the lenght of the side and therefore they will change at the same rate

OpenStudy (anonymous):

i got dz/dt = 2√3

OpenStudy (anonymous):

2* radical 3

OpenStudy (anonymous):

oh sorry, yes of course it wil be \(\sqrt{3} \times\) times the rate the edges change You're right